OpenStudy (anonymous):

Rationalize the denominator:

OpenStudy (anonymous):

\[\frac{ 2\sqrt{3} -5}{ \sqrt{3} +4}\]

OpenStudy (anonymous):

\[\frac{2\sqrt3 - 5} {\sqrt3 +4} = \frac{(2\sqrt3 - 5)(\sqrt3 - 4)}{3 - 16}\]

OpenStudy (anonymous):

Oh... thanks